Thursday, September 19, 2013

Blue Balls Be Back!

The custom screw-ball tools had cured sufficiently over night so the first job of the day was the ball replacement surgery for the blue camera...

The anaesthesiologist was at the ready.

The patient was prepped...

Then (apologies for the mixed metaphors) Lights, Camera & Action.

The steady hands of Chief Surgeon Bramall.  The micrometer was read & then unlocked & backed off to make space to insert the ball.

OK: 1 ball back in place & the screw snapped off!  Residual glue to be removed later.

Now for the second one...

Longnose pliers provided the extra reach required to get the tool + ball into position.

In it goes, visibility aided by having the upper hatch open as well.

Here's the ball + tool in position, before breaking the glue to remove the bolt.

The next trick was getting rid of the glue that mostly stayed behind on the ball.  Scalpel please nurse...

Right - there's #2, back where it belongs.  Then reset the micrometers to their pre-surgery positions.

David kindy offered me the chance to poke around inside the dewar to get a feel for how tight the springs are.  Hopefully it's not something I really need to know, but it's all part of getting acquainted with our amazing new toy :)

After all that, the patient was sewn back up & connected to the vacuum pump.  Once a reasonable vacuum's achieved, we can switch on the cooling to get the chip down to its operational tempearture of -110 °C.  Only then will we be able to tell whether the procedure's eliminated the focus gradient.

Otherwise - Eben had machined down the little pieces of G10 that hold the exposure meter mirrors in place so these could be installed.

& Jürgen also installed the calibration fibre next to the ThAr lamp on the bench outside the tank.

By then it was time for the afternoon tour - David Buckley joining us this time for an audience with Her Royal Spectrographness.

Driving down the hill for dinner we could see some heavy weather coming our way... 

We returned later in the evening to see how the cool-down was progressing. 

& took a couple of frames with a halogen lamp on to make sure the chip had survived its morning ordeal.  Happily it had & we could see that the focus was way off.  

This was a good sign since we'd previously moved the pupil mirror to try to achieve focus with the CCD badly tilted.  Now with the chip (hopefully!) back where it should be - the pupil mirror could go back to where it belongs.

We'll see tomorrow, once the system's properly cooled, whether we're sorted or not...

No comments:

Post a Comment